A software product, consisting of 12,200 source lines of code, was developed by four programmers, each working 20 hours per week, over a six-month period. After delivery, approximately 160 source lines are modified every month. What is the annual maintenance effort for this product in person-months

Here is the equation:

AME = ACT X SDT, where

AME is the annual maintenance effort measured in person months
ACT is the annual change traffic, which represents a fraction of a software product's source instructions that undergo change during a typical year through addition or modification
SDT is the software development time in person months

The final answer should be 1.884, but how can I get that?
Answer #1

Since AMC is driven by the product of ACT and SDT , calculating the 2 values as shown below:

ACT or annual change traffic is given by fraction of the source code that gets changed in a typical year

which here is 160 lines per month out of (12,200/6) lines per month

Annually, the fraction would be (160*12)/(12,200/6*12)= 0.078688

Now, let us calculate SDT that is software development time in person months

4 developers take 6 months to complete a project which happens to be straightforward 24 person-months of effort

Now, AMC = ACT * SDT

= 0.078688 * 24

= 1.88